# GPU Memory and Utilization Estimation for Training-Aware Resource Management: Opportunities and Limitations

### v1

@@ 4.1.1. Data Cleaning and Feature Extraction
For each architecture type, a specialized data cleaner script parses the corresponding logs and summary files. These scripts extract:
@@
  • System metrics (e.g., CPU%, MEM%)
The cleaned and structured datasets are saved in a standardized format for seamless downstream processing. While the primary targets are GPU memory usage and utilization metrics (SMACT, SMOCC, DRAMA), the pipeline preserves a broad set of features to support future research directions, such as energy consumption or latency prediction. Data cleaning scripts are configured by specifying the raw dataset directory and the desired output CSV path.
4.1.2. Analysis and Modeling
Subsequent analysis is performed through detailed Jupyter Notebooks, included in the repository. These notebooks:
  • Visualize the collected data and distributions
  • Provide a step-by-step guide for training memory usage estimators using MLP-based and Transformer-based models
  • Include preprocessing, training, and evaluation logic
4.1.3. Estimator Training and Evaluation
We implement and evaluate ensemble-based estimators for memory usage prediction. Each ensemble aggregates multiple base models with varied hyperparameters and structures to improve generalization. Models are trained and evaluated using 3-fold cross-validation, ensuring robust performance reporting across all architecture families.
4.1.4. Testing on unseen real-world models
The repository provides parsers that extract input features from real-world models and feed them into the trained memory estimator model. The inference for normal and ensemble model are different as their architecture is different.
4.1.5. Considerations for Collecting the data
We follow these design considerations both in the data collection scripts and when configuring the launcher script.
Architecture over model types. To ensure long-term relevance, we focus on architecture families (e.g., CNNs, Transformers) rather than specific models, which quickly become outdated.
4.2. Classification vs. Regression.
GPU memory usage exhibits a staircase-like growth pattern (Figure 7), observed in our experiments, making regression challenging due to sharp discontinuities, flat plateaus with weak gradients, and spiking errors at step boundaries. These issues lead to instability and overfitting in regression models.
@@ 4.1.5. Considerations for Collecting the data
We formulate the estimation task as classification by discretizing memory usage into fixed-size bins (e.g., 0–1,GB, 1–2,GB). Since CNNs and Transformers exhibit significantly higher memory demands than MLPs, larger bin sizes are used for these architectures to better capture their distribution. Regression models yielded poor performance for both memory and utilization metrics, motivating the use of classification throughout. For GPU utilization metrics, class imbalance and limited coverage of the metric range posed additional challenges, requiring bin adjustments. PCA visualization in Figure 4 further reveals that utilization metrics lack clear class separation. This stands in contrast to GPU memory, where more distinct class boundaries are observed—specifically for MLPs—whereas the borders for CNNs and Transformers remain ambiguous, causing the model to become confused when learning these patterns.
4.3. Selected Features
@@
For all datasets, the MLP-based estimator uses: batch size, number of parameters, number of activations, #activations $\times$ batch size and the number of architecture-specific layer types (e.g., linear, convolutional, batch normalization, and dropout layers). The MLP and CNN datasets include activation encoding (cos/sin). The Transformer-based estimator extends all of the above with a sequence of (layer type, #activations, #parameters) tuples. These features capture model scale and structure — the primary determinants of GPU memory requirements and utilization.
@@
Table 6 in Appendix presents the target variable ranges alongside total parameters and activations extracted from model summaries. GPU Memory estimator test models are trained with cross-entropy loss (Adam optimizer) and evaluated via stratified 3-fold cross-validation (70% train / 30% validation per fold, with a held-out 30% test set), reporting Accuracy and F1-score to measure overall correctness and class balance, respectively.
@@
5. Methodology
6. Evaluation
Results for the GPU Memory. GPUMemNet achieves high accuracy on MLPs (serving as a proof-of-concept) and strong performance on CNNs and Transformers (0.81–0.88 at 8 GB memory range bins), supporting the classification formulation for practical memory estimation (Table 2). Using an 8 GB bin width preserves accuracy but can limit collocation opportunities, so we experimented with smaller bins. For CNNs and Transformers, however, finer bins degraded performance—reflecting complex, high-variance patterns that would require substantially more data to learn, incurring significant time and hardware costs (a core caveat of data-driven estimation). A second challenge is drift: any change in frameworks or memory optimizations can invalidate labels, forcing full data recollection and reprocessing. Third, generalization is brittle: ML estimators extrapolate poorly to unseen architectures or new layers. For example, probing a GPT-2–style model revealed 1D convolution layers—operators absent from GPUMemNet’s Transformer training set—explaining poor extrapolation. This underscores a practical need: keep the GPUMemNet datasets open and continuously extended to cover new layers and variants, improving accuracy over time. Finally, the runtime cost of a GPUMemNet estimator model is low: at most 16 ms on an NVIDIA A100 (40 GB) and 32 ms on an AMD EPYC CPU, measured over 100 runs per estimator (Table 2).
Results for the GPU utilization. Table 7 in Appendix shows clear performance disparities across architectures. Accuracy drops significantly for non-MLP architectures, with DRAMA scoring as low as 62% on Transformers. Recall — critical for avoiding false low-utilization classifications — is similarly poor across most models. As shown in Figure 6 and discussed in Section 4.2, utilization classes lack clear separation, with CNN and Transformer data points heavily overlapping, hindering meaningful decision boundaries. Furthermore, single-workload utilization prediction assumes additivity, which does not hold due to non-linear inter-workload interactions across execution phases.
Memory Estimators on Diverse Real-World Models. After the in-depth look at the individual strengths and weaknesses of the different GPU memory need estimation methods, we compare all three estimators—the Horus formula (Yeung et al., 2022), FakeTensor (Contributors, 2025), and GPUMemNet—on real models in Figure 5. We use GPUMemNet’s MLP-based estimators throughout, given their higher accuracy on CNNs and Transformers (Table 2). The Horus formula can both under- or overestimate—underestimates risk OOM, while overestimates waste memory and reduce collocation. FakeTensor typically underestimates (risking OOM) and, for many Transformer models, produces no estimate at all due to unsupported ops. In contrast, GPUMemNet tends to overestimate memory, reducing potential collocation gains. Its largest error appears on GPT-2—an out-of-distribution case for our training data—where unseen architectural elements drive the mismatch. Furthermore, it cannot estimate for the DLRM as it has not been trained for it, underlining the weakness of model-based estimation on unseen cases.
Utilization Estimator on Diverse Real-World Models.
@@ 7. Discussion and Limitations
Overcoming Limitations of Existing Estimators. Current estimation tools often suffer from intrusive designs that require deep integration into upper software layers, making them impractical for general use. Furthermore, many existing solutions struggle with high latency overhead on the scheduler’s critical path, frequent misestimations, and a general inability to generalize across different models and GPU model heterogeneity.
@@
The Challenges of Estimating GPU Utilization. GPU utilization estimation remains an understudied and challenging area where our initial attempts were not successful. Using utilization as a proxy for resource interference is controversial because these metrics are not additive and vary across GPU models, suggesting that future work must address these non-linear hardware complexities.
@@ 8. Conclusion
In this paper, we investigated GPU memory and utilization estimation, providing a comprehensive qualitative and quantitative analysis of their respective trade-offs. We presented a data-driven, extensible framework designed to estimate resource requirements for deep learning training tasks. By addressing the practical limitations of current estimation methods, more informed and efficient resource management within GPU clusters can be achieved.

### v2

@@ 3.1.3. Data Cleaning and Feature Extraction
For each model architecture type, a specialized data cleaner script parses the corresponding logs and summary files. These scripts extract:
@@
  • CPU metrics (e.g., CPU%, MEM%)
The cleaned datasets are saved in CSV format for downstream processing. While the primary metrics in this work are GPU memory usage and utilization (SMACT, SMOCC, DRAMA) (Yousefzadeh-Asl-Miandoab et al., 2023), the dataset includes broader set of features to support future research directions, such as energy consumption or latency prediction.
3.2. Classification vs. Regression
As shown in Figure 2, GPU memory usage exhibits a staircase-like growth pattern. This makes approaching the GPU memory use estimation as a regression challenging due to sharp discontinuities, flat plateaus with weak gradients, and spiking errors at step boundaries. These issues lead to instability and overfitting in regression models.
@@
Therefore, we formulate the estimation task as classification by discretizing memory usage into fixed-size bins (e.g., 0–1,GB, 1–2,GB). Since CNNs and Transformers exhibit significantly higher memory demands than MLPs, larger bin sizes are used for these architectures to better capture their distribution.
@@
For estimation of the GPU utilization metrics, class imbalance and limited coverage of the metric range pose additional challenges, requiring bin adjustments. PCA visualization in Figure 6 further reveals that utilization metrics lack clear class separation. This stands in contrast to GPU memory, where more distinct class boundaries are observed in Figure 3—specifically for MLPs—whereas the borders for CNNs and Transformers remain ambiguous, causing the model to become confused when learning these patterns.
3.3. GPU Memory Need Estimation
We implement and evaluate ensemble-based estimators for memory usage prediction. For the estimator models, we evaluate two types of ensemble models for better accuracy and stability during training: an ensemble of MLPs (Figure 7(a) in Appendix) and Transformer classifiers (Figure 7(b) in Appendix).
Each ensemble aggregates multiple base models with varied hyperparameters and structures to improve generalization. Models are trained and evaluated using 3-fold cross-validation, ensuring robust performance reporting across all architecture families. We test the estimator using different deep learning model architectures as well, where the estimators can be MLP- and Transformers-based.
@@ 3.3. GPU Memory Need Estimation
For all datasets (Section 3.1.2), the MLP-based estimator uses: batch size, number of parameters, number of activations, #activations $\times$ batch size and the number of architecture-specific layer types (e.g., linear, convolutional, batch normalization, and dropout layers). The MLP and CNN datasets include activation encoding (cos/sin), which is also added to the set of features for estimator training. The Transformer-based estimator extends all of the above with a sequence of (layer type, #activations, #parameters) tuples. These features capture model scale and structure — the primary determinants of GPU memory requirements and utilization.
@@
Table 6 in Appendix presents the target variable ranges alongside total parameters and activations extracted from model summaries. GPU Memory estimator, GPUMemNet, test models are trained with cross-entropy loss (Adam optimizer) and evaluated via stratified 3-fold cross-validation (70% train / 30% validation per fold, with a held-out 30% test set), reporting Accuracy and F1-score to measure overall correctness and class balance, respectively.
3.4. GPU Utilization Estimator
Given that GPU utilization metrics remain largely understudied compared to GPU memory estimation and the Horus dataset and models are currently unavailable, we train a deep learning-based estimator to predict three utilization metrics — SMACT, SMOCC, and DRAMA — using the dataset we curated for GPU memory estimation. These metrics, averaged over the execution lifetime of each training job, together capture how both the compute and memory units of the GPU are utilized, providing a more complete picture of resource usage beyond memory consumption alone. Following the same methodology applied to GPU memory estimation, we reuse the same input features and dataset, applying the same empirical evaluation approach to assess their predictive power.
4. Quantitative Analysis of Estimators
4.1. Experimental Environment
4.2. Evaluated Estimators
From the three classes of GPU memory need estimators surveyed in Section 2.1, we select one available method per paradigm for empirical evaluation, the Horus analytical formula (Yeung et al., 2022), PyTorch FakeTensor (Contributors, 2025), and our own lightweight ML-based estimator, GPUMemNet. The reasons for not including others are DNNMem (Gao et al., 2020), which is not publicly available, LLMem (Kim et al., 2024), which is restricted to transformer decoder fine-tuning, and DeepSpeed’s estimator (Microsoft DeepSpeed Team, 2023), which is scoped exclusively to ZeRO-stage configurations and does not account for activation memory or framework overhead. DNNPerf (Gao et al., 2023) relies on a GNN operating over the model’s computation graph, which requires framework deserialization APIs to parse and reconstruct operator-level graph representations — a more complex and tightly coupled integration compared to a lightweight feature-based approach. Furthermore, its GNN-based inference latency scales with model and graph complexity, and constructing sufficiently representative training datasets for GNN-based estimators carries significant collection and curation overhead, motivating the design of our own lightweight estimator, GPUMemNet.
4.3. Memory Estimators on Micro-benchmarks
Analytical. Horus (Yeung et al., 2022) estimates training memory via a lightweight analytical pass over model statistics, incurring negligible overhead. However, as Figure 4 demonstrates, it systematically overestimates memory consumption across MLP configurations due to its inability to account for framework-level optimizations such as activation reuse, dynamic memory management, and layer fusion. While this conservatism avoids OOM failures, it unnecessarily reduces collocation opportunities by over-reserving memory.
CPU-side Libraries, tools. PyTorch FakeTensor (Contributors, 2025) performs shape propagation — optionally symbolic — creating lightweight tensor surrogates that carry metadata (shape, dtype, device, and aliasing information) without allocating real memory, enabling analysis of model structure without modifying or running the training process itself. Evaluated across 2,030 model configurations from the TIMM (Wightman, 2019) library, FakeTensor consistently underestimates peak memory usage (Table 1), as it does not account for the caching allocator’s reserved memory, intermediate buffers, or optimizer states. A fixed safety margin (e.g., +4,GB) can partially compensate for this, providing a conservative bound suitable for collocation decisions. Its runtime overhead, while modest, is the most prominent among the three candidates and adds latency to the scheduling critical path. Additionally, FakeTensor fails on architectures that rely on unsupported operations, dynamic control flow, or CUDA-only kernels — including many Transformer variants, DLRM, and Mask R-CNN — limiting its practical coverage.
P50 P80 P90 P95 P99 Max
Absolute error (GB) 1.16 1.97 2.82 3.84 6.82 (>> 8GB)
Estimation time (s) 0.94 1.49 1.84 2.28 3.47 4.97
ML-based. GPUMemNet (Section 3), is trained on a curated synthetic dataset spanning MLP, CNN, and Transformer architecture families and formulates memory prediction as a classification task over discretized memory ranges. Unlike the analytical and library-based alternatives, it requires no model source code, no architectural specification beyond high-level configuration parameters, and introduces minimal inference latency.
Results in Table 2 show that GPUMemNet achieves high accuracy on MLPs (serving as a proof-of-concept) and strong performance on CNNs and Transformers (0.81–0.88 at 8 GB memory range bins), supporting the classification formulation for practical memory estimation. Using an 8 GB bin width preserves accuracy but can limit collocation opportunities, so we experimented with smaller bins. For CNNs and Transformers, however, finer bins degraded performance—reflecting complex, high-variance patterns that would require substantially more data to learn, incurring significant time and hardware costs (a core caveat of data-driven estimation). A second challenge is drift: any change in frameworks or memory optimizations can invalidate labels, forcing full data recollection and reprocessing. Third, generalization is brittle: ML estimators extrapolate poorly to unseen architectures or new layers. For example, probing a GPT-2–style model revealed 1D convolution layers—operators absent from GPUMemNet’s Transformer training set—explaining poor extrapolation. This underscores a practical need: keep the GPUMemNet datasets open and continuously extended to cover new layers and variants, improving accuracy over time. Finally, the runtime cost of a GPUMemNet estimator model is low: at most 16 ms on an NVIDIA A100 (40 GB) and 32 ms on an AMD EPYC CPU, measured over 100 runs per estimator (Table 2).
@@ 4.3. Memory Estimators on Micro-benchmarks
4.4. Memory Estimators on Real-World Models
After the in-depth look at the individual strengths and weaknesses of the different GPU memory need estimation methods, we compare all three estimators—the Horus formula (Yeung et al., 2022), FakeTensor (Contributors, 2025), and GPUMemNet—on diverse real-world models in Figure 5. We use GPUMemNet’s MLP-based estimators throughout, given their higher accuracy on CNNs and Transformers (Table 2). The Horus formula can both under- or overestimate—underestimates risk OOM, while overestimates waste memory and reduce collocation. FakeTensor typically underestimates (risking OOM) and, for many Transformer models, produces no estimate at all due to unsupported ops. In contrast, GPUMemNet tends to overestimate memory, reducing potential collocation gains. Its largest error appears on GPT-2—an out-of-distribution case for our training data—where unseen architectural elements drive the mismatch. Furthermore, it cannot estimate for the DLRM as it has not been trained for it, underlining the weakness of model-based estimation on unseen cases.
4.5. Utilization Estimator
@@ 5. Discussion and Limitations
Overcoming Limitations of Existing Estimators. Current estimation tools often suffer from intrusive designs that require deep integration into upper software layers, making them impractical for general use. Furthermore, many existing solutions struggle with high latency overhead on the scheduler’s critical path, frequent misestimations, and a general inability to generalize across different models and GPU generations.
@@
The Challenges of Estimating GPU Utilization. GPU utilization estimation remains an understudied and challenging area where our initial attempts were not successful. Using utilization as a proxy for resource interference is not straightforward because these metrics are not additive and vary across GPU architectures, suggesting that future work must address these non-linear hardware utilization characteristics.
@@ 6. Conclusion
In this paper, we investigated GPU memory and utilization estimation, providing aqualitative and quantitative analysis of their respective trade-offs. We presented a data-driven, extensible framework designed to estimate resource requirements for deep learning training tasks. By addressing the practical limitations of current estimation methods, more informed and efficient resource management within GPU clusters can be achieved.